# LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation

### v1

@@ 4.3 LogitSpec Verification
After retrieving multiple draft token sequences $\mathcal{D}$, we organize these token sequences into a draft tree and prepare a tree attention for parallel verification. Intuitively, this process involves flattening $\mathcal{D}$ into a draft token sequence $\mathbf{x}_{\text{draft}}$.
Then, we prepare an attention mask to make each draft token sequence invisible to other sequences. Let $\text{Tril}(l)$ denotes a causal mask with length $l$, and $m_{j}$ denotes the length of the $j$-th token sequence in $\mathcal{D}$, the attention mask is given by:
$A_{\text{draft}}=\operatorname{diag}(\text{Tril}(m_{1}),\text{Tril}(m_{2}),...,\text{Tril}(m_{j})).$ (6)
@@ 5.1 Experimental Setup
We conduct experiments on various text generation benchmarks to evaluate the effectiveness of LogitSpec. Specifically, we first evaluate the performance of LogitSpec on Spec-Bench (Xia et al., 2024). Spec-Bench is a widely used comprehensive benchmark that covers diverse application scenarios, including multi-turn conversation (MT), translation (Trans), summarization (Sum), question answering (QA), mathematical reasoning (Math), retrieval-augmented generation (RAG). Furthermore, we evaluate the performance of LogitSpec on HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021) and CNN/DM (Nallapati et al., 2016), which are widely used text generation benchmarks in existing works.
Baselines.
In our main experiments, we compare our LogitSpec to existing state-of-the-art plug-and-play SD methods: (a) Lookahead Decoding (Fu et al., 2024): utilizing the dropped draft tokens to serve as the retrieval model; (b) REST (He et al., 2024): employing an external knowledge base as the retrieval model; (c) PLD (Saxena, 2023): extracting n-grams from the user-input prompts and decoded tokens as the retrieval model; and (d) vanilla SpS (Leviathan et al., 2023; Chen et al., 2023): employing a well trained small-size model Vicuna-68M as the draft model.
@@
For all experiments, we follow the standard setting of Spec-Bench with Pytorch (Paszke et al., 2019) version of 2.6.0, transformers (Wolf et al., 2020) version of 4.37.1, and CUDA (NVIDIA et al., 2020) 12.4. The experiments where the model size is between 7B and 33B are conducted on single NVIDIA SXM A100 80G GPU, while the experiments for models with 70B size are conducted on 2 NVIDIA SXM A100 80G GPUs. The inference precision is set as float16. Following prior work, we conduct experiments with temperature 0 with batch size of 1. The maximal generation lengths are set to 1024. More implementation details can be found in Appendix D. In our experiments, we use two widely used metrics for evaluation: mean accepted tokens per decoding step (denoted as MAT) and overall speedup ratio (denoted as Speedup).
@@ 5.2 Main Results
We conduct extensive experiments on various text generation benchmarks to demonstrate the effectiveness of our LogitSpec. As presented in Tables 1 and 2, LogitSpec significantly outperforms retrieval-based SD baselines by a large margin. We further present the experiments of Llama 2 chat series models in Appendix E.
@@
Models Method CNN/DM GSM8K HumanEval Average
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Vicuna 7B Lookahead 1.54 1.28 1.92 1.64 1.79 1.57 1.75 1.50
REST 1.64 1.19 1.55 1.13 1.96 1.51 1.72 1.28
PLD 2.61 2.26 1.80 1.62 1.84 1.65 2.08 1.84
SpS 2.34 1.58 2.08 1.49 2.56 1.80 2.33 1.62
LogitSpec 3.28 2.61 2.69 2.20 2.62 2.24 2.86 2.35
Vicuna 13B Lookahead 1.46 1.12 1.88 1.61 1.75 1.57 1.70 1.43
REST 1.65 1.22 1.57 1.18 1.94 1.55 1.72 1.32
PLD 2.34 1.85 1.76 1.61 1.98 1.77 2.03 1.74
SpS 2.18 1.48 2.00 1.55 2.66 1.95 2.28 1.66
LogitSpec 2.90 2.17 2.59 2.13 2.88 2.47 2.79 2.26
Vicuna 33B Lookahead 1.50 1.18 1.89 1.44 1.66 1.33 1.68 1.32
REST 1.65 1.41 1.57 1.32 1.99 1.64 1.74 1.46
PLD 2.07 1.72 1.66 1.41 1.60 1.42 1.78 1.52
SpS 2.01 1.57 1.97 1.54 2.27 1.80 2.08 1.64
LogitSpec 2.66 2.01 2.46 1.93 2.41 1.95 2.51 1.96
Models Method MT Trans Sum QA Math RAG MAT Speedup
Vicuna 7B Lookahead 1.40 1.14 1.19 1.24 1.55 1.09 1.66 1.27
REST 1.63 1.31 1.36 1.66 1.21 1.73 1.82 1.48
PLD 1.64 1.04 2.43 1.14 1.61 1.71 1.73 1.59
SpS 1.66 1.13 1.62 1.49 1.47 1.55 2.28 1.49
LogitSpec 1.92 1.39 2.68 1.70 2.22 1.87 2.44 2.01
Vicuna 13B Lookahead 1.30 1.06 1.20 1.12 1.48 1.12 1.63 1.22
REST 1.52 1.17 1.37 1.53 1.19 1.55 1.82 1.38
PLD 1.47 1.02 2.19 1.03 1.57 1.71 1.68 1.48
SpS 1.60 1.13 1.68 1.39 1.53 1.67 2.18 1.49
LogitSpec 1.89 1.43 2.33 1.43 2.23 1.93 2.32 1.93
Vicuna 33B Lookahead 1.32 1.08 1.20 1.06 1.54 1.15 1.61 1.24
REST 1.63 1.27 1.45 1.61 1.30 1.61 1.80 1.48
PLD 1.44 1.06 2.00 1.07 1.55 1.45 1.55 1.42
SpS 1.75 1.28 1.76 1.53 1.69 1.68 2.01 1.61
LogitSpec 1.77 1.38 2.15 1.37 2.00 1.69 2.13 1.76
@@ 5.3 Ablation Study
To further investigate the components of LogitSpec and provide more insights, we conduct extensive ablation studies on the aforementioned 4 benchmarks in Table 3. Specifically, we mainly focus on two components of LogitSpec: last logit and retrieval. We denote LogitSpec without the last logit decoding and only using retrieving as w/o last logit, and LogitSpec without retrieval and only using last logit decoding as w/o retrieval. As shown in Table 3, the absence of any component results in a performance degradation of the entire framework.
Our findings are as follows First, the absence of retrieval exhibits more importance to the final acceleration, which is consistent with the discussion in Section 4.2 that the theoretical upper bound of last logit decoding severely hinders its real-world application. Second, the absence of retrieval and the absence of last logit shows different affect in different sub-tasks. For example, the absence of retrieval decreases MAT of 0.87 on Spec-Bench with Vicuna 7B, while it decreases MAT of 1.71 on CNN/DM. In contrast, the absence of last logit leads to a more consistent MAT degradation across different tasks, which further highlights the robustness of last logit decoding. Finally, these results demonstrate that combining last logit with retrieval improve the retrieval performance and overall speedup.
@@
5.4 Case Study
In-depth Running Time Analysis.
To further investigate the effectiveness of LogitSpec, we conduct experiments to analyze the running time allocation within the whole decoding process. Specifically, there are five non-negligible components in LogitSpec, including (a)retrieving draft tokens: the process of retrieving reference for the next token and the next next tokens; (b) preparation: preparing attention mask for the draft tokens; (c) model forward: conducting one-pass model forward; (d) verification: validating the draft tokens with speculative sampling; (e) update: necessary update of KV cache and retrieval model. We present the running time breakdown in Figure 4.
As shown in Figure 4, model forward occupies the majority of wall-clock time. Compared with vanilla auto-regressive decoding, the overhead introduced by LogitSpec, i.e., retrieval overhead, only takes 1.17% of the whole decoding process. This overhead can be further alleviated by some parallel techniques, as the retrieval process of the next next token is independent. The process of verification and update takes 1.03% and 1.05% through the whole process, respectively, which brings negligible overhead as well.
Retrieval Performance.
As mentioned in Section 4.2, our LogitSpec can effectively expand the retrieval range and improve the retrieval performance. To validate this statement, we conduct a simple experiment to test the retrieval successful rate, i.e., whether the model successfully retrievals reference as draft tokens. As shown in Table 4, PLD cannot retrieval any matched tokens in more than 30% decoding steps, while LogitSpec can retrieval matched reference in most decoding steps. These results further demonstrate the effectiveness of LogitSpec.
@@ 6 Conclusion
Conclusion.
[…]
Limitations and ethic statements.
While our LogitSpec is a fully plug-and-play SD framework, it is possible to integrate a draft model for further acceleration. We affirm our commitment to contributing positively to society, avoiding harm, and upholding honesty and trustworthiness. We appropriately cite the previous methods and datasets we use, and ensure that all data involved is fully public, with no private data being utilized.
@@ Appendix A More Discussions to Related Work
We give more discussions to existing works in both draft-model-based speculative decoding and draft-model-free speculative decoding.
@@
Besides the discussions in Section 2, there are also many excellent works related to the draft-model-based speculative decoding. For example, HASS [Zhang et al., 2025] discovers the inconsistency of EAGLE between training and inference, and proposes a multi-step training framework to address this. CORAL [Weng et al., 2025] proposes a cross-step representation alignment to address this problem. Judge Decoding [Bachmann et al., 2025] recognizes the potential of accepting high-quality but refused draft tokens to further improve the acceleration. Gumiho [Li et al., 2025c] demonstrates that the initial draft token is more important and proposes a hybrid model to combine serial and parallel draft head.
Besides these methods that focus on the training process, the process of verification also draws extensive interests, mainly focus on adaptive draft length. SpecDec++ [Huang et al., 2025] formulates the verification process as a Markov decision process to adaptively determine the draft length. OPT-Tree [Wang et al., 2024] proposes a method to search the optimal tree structure that maximizes the mathematical expectation of the acceptance length in each decoding step. AdaEAGLE [Zhang et al., 2024b] proposes a novel framework to explicitly model the draft tree structure for EAGLE series models. PEARL [Liu et al., 2025] pioneers this direction by serving the draft model and the target model in parallel to achieve a segmented adaptive draft length.
@@
For layer sparsity, SPEED [Hooper et al., 2024] proposes a method to speculatively execute multiple future tokens in parallel with the current token using predicted values based on early-layer hidden states. FREE [Bae et al., 2023] proposes a shallow-deep module and a synchronized parallel decoding to improve the efficiency. EESD [Liu et al., 2024] proposes an early-exiting framework with a self-distillation method and leverages Thompson Sampling to regulate the generation processes. For retrieval-based SD, Token Recycling [Luo et al., 2024] proposes a method to utilize the dropped draft tokens and generate draft tokens via an adjacency matrix. Different with our method, Token Recycling is a query-dependent method that utilize information from other queries, which may result in limitations when applied to real-world applications with complex and dynamic user inputs. SAM Decoding [Hu et al., 2024] utilizes common text corpus and dynamic text sequence as retrieved sources and proposes suffix automaton to efficiently obtain yields more accurate match positions.
Appendix B Experimental Setup for the Motivation Experiments
To investigate the prediction ability of the last logit, we conduct two motivation experiments to demonstrate the effectiveness and robustness of the last logit. For the effectiveness of the last logit, as shown in Figure 2, we conduct auto-regressive inference for the 6 models and record the logits for each decoded token. Then, for each decoded token $x_{i}$ which is sampled from the logit $p_{i-1}$, we investigate the rank of $x_{i}$ in $p_{i-2}$, i.e., the corresponding last logit, and visualize the statistics of the rank in Figure 2. For the robustness of the last logit, as shown in Figure 2, we conduct last logit decoding and investigate its mean accepted tokens per decoding step. Both the two experiments are conducted on a small subset of Spec-Bench, where we randomly sample 2 questions for each sub-category of MT, and 10 questions for other categories (Trans, Sum, QA, Math, RAG).

### v2

@@ 4.3 LogitSpec Verification
After retrieving multiple draft token sequences $\mathcal{D}$, we organize these token sequences into a draft tree and prepare a tree attention for parallel verification.
Then, we prepare an attention mask to make each draft token sequence invisible to other sequences. Let $\Lambda(l)$ denote a causal mask with length $l$, and $m_{j}$ denote the length of the $j$-th token sequence in $\mathcal{D}$; the attention mask is given by:
$A_{\text{draft}}=\operatorname{diag}\big(\Lambda(m_{1}),\ldots,\Lambda(m_{j})\big).$ (8)
@@ 5.1 Experimental Setup
We evaluate LogitSpec on a broad suite of text generation benchmarks. We first use Spec-Bench (Xia et al., 2024), a widely used comprehensive benchmark that covers diverse application scenarios, including multi-turn conversation (MT), translation (Trans), summarization (Sum), question answering (QA), mathematical reasoning (Math), and retrieval-augmented generation (RAG). We then further evaluate LogitSpec on HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021), CNN/DM (Nallapati et al., 2016), MATH (Hendrycks et al., 2021), AIME 24&25 (Mathematical Association of America, 2025), and LongBench (Bai et al., 2024), which are widely used benchmarks for code generation, math reasoning, summarization, and long-context generation.
@@
We follow standard inference settings with temperature 0 and batch size 1. All implementation details (baselines, hardware, software versions, model configurations) are provided in Appendix D. We use two widely used metrics for evaluation: mean accepted tokens per decoding step (denoted as MAT) and overall speedup ratio (denoted as Speedup). We further evaluate LogitSpec under non-zero temperatures ($T>0$) in Appendix E.4.
@@ 5.2 Main Results
We conduct experiments on various text generation benchmarks to demonstrate the effectiveness of LogitSpec. As in Table 1 (with the Spec-Bench breakdown deferred to Appendix E.2), LogitSpec outperforms retrieval-based SD baselines by a large margin. We further present more in-depth analysis of LogitSpec including (i) more LLM backbones (Llama2 series, Llama-3.1-Instruct-8B and Qwen-3-8B), (ii) more benchmarks (MATH and AIME datasets), and (iii) long context benchmarks (the LongBench dataset) in Appendix E.
@@
Models Method CNN/DM GSM8K HumanEval Spec-Bench Average
MAT Speedup MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Vicuna 7B Lookahead 1.54 1.28 1.92 1.64 1.79 1.57 1.66 1.27 1.73 1.44
REST 1.64 1.19 1.55 1.13 1.96 1.51 1.82 1.48 1.74 1.33
PLD 2.61 2.26 1.80 1.62 1.84 1.65 1.73 1.59 2.00 1.78
SpS 2.34 1.58 2.08 1.49 2.56 1.80 2.28 1.49 2.32 1.59
LogitSpec 3.28 2.61 2.69 2.20 2.62 2.24 2.44 2.01 2.76 2.26
Vicuna 13B Lookahead 1.46 1.12 1.88 1.61 1.75 1.57 1.63 1.22 1.68 1.38
REST 1.65 1.22 1.57 1.18 1.94 1.55 1.82 1.38 1.75 1.33
PLD 2.34 1.85 1.76 1.61 1.98 1.77 1.68 1.48 1.94 1.68
SpS 2.18 1.48 2.00 1.55 2.66 1.95 2.18 1.49 2.26 1.62
LogitSpec 2.90 2.17 2.59 2.13 2.88 2.47 2.32 1.93 2.67 2.18
Vicuna 33B Lookahead 1.50 1.18 1.89 1.44 1.66 1.33 1.61 1.24 1.67 1.30
REST 1.65 1.41 1.57 1.32 1.99 1.64 1.80 1.48 1.75 1.46
PLD 2.07 1.72 1.66 1.41 1.60 1.42 1.55 1.42 1.72 1.49
SpS 2.01 1.57 1.97 1.54 2.27 1.80 2.01 1.61 2.07 1.63
LogitSpec 2.66 2.01 2.46 1.93 2.41 1.95 2.13 1.76 2.42 1.92
@@ 5.3 Ablation Study
To further provide more insights of LogitSpec, we conduct extensive ablation studies on the aforementioned 4 benchmarks in Table 2. Specifically, we mainly focus on two components of LogitSpec: last logit and retrieval. We denote LogitSpec without the last logit decoding and only using retrieval as w/o last logit, and LogitSpec without retrieval and only using last logit decoding as w/o retrieval. As shown in Table 2, the absence of any component results in a performance degradation of the entire framework.
Our findings are as follows. First, the absence of retrieval exhibits more importance to the final acceleration, which is consistent with the discussion in Section 4.2 that the theoretical upper bound of last logit decoding severely hinders its real-world application. Second, the absence of retrieval and the absence of last logit show different effects in different sub-tasks. For example, the absence of retrieval decreases MAT by 0.87 on Spec-Bench with Vicuna 7B, while it decreases MAT by 1.71 on CNN/DM. In contrast, the absence of last logit leads to a more consistent MAT degradation across different tasks, which further highlights the robustness of last logit decoding. Finally, these results demonstrate that combining last logit with retrieval improves the retrieval performance and overall speedup. We also conduct ablation studies on pruning strategies in Appendix E.7 and E.8. The experimental results suggest that varying the pruning strategy yields only minor differences.
@@
5.4 Case Study
In-depth Running Time Analysis.
We conduct experiments to analyze the running time allocation of the whole decoding. Specifically, there are five non-negligible components in LogitSpec, including (a) retrieving draft tokens: the process of retrieving reference for the next token and the next next tokens; (b) preparation: preparing attention mask for the draft tokens; (c) model forward: conducting one-pass model forward; (d) verification: validating the draft tokens with speculative sampling; (e) update: necessary update of KV cache and retrieval model.
As shown in Figure 5, model forward occupies the majority of wall-clock time. Compared with vanilla AR decoding, the overhead introduced by LogitSpec, i.e., retrieving overhead, only takes 1.17% of the whole decoding process. It can be further alleviated by parallel techniques, as the retrieval process is independent, which brings negligible overhead as well.
Retrieval Performance.
As mentioned in Section 4.2, LogitSpec expands the retrieval range and improves the retrieval performance. We conduct a simple experiment to test the retrieval success rate, i.e., whether the model successfully retrieves the reference as draft tokens. As in Table 3, PLD cannot retrieve any matched tokens in more than 30% decoding steps, while LogitSpec retrieves matched reference in most decoding steps. These results further demonstrate the effectiveness of LogitSpec.
Real-world examples.
We also provide a real-world example in Appendix E.9 to illustrate how next-next token speculation allows LogitSpec to succeed where standard retrieval methods fail.
@@ 6 Conclusion
In this paper, we observe that the logit of the last token can predict the next next token with a relatively high accuracy without any fine-tuning. Based upon this observation, we propose a novel retrieval-based SD framework, namely LogitSpec, which utilizes the prediction ability of the last logit to effectively expand the retrieval range and find the most relevant reference as the draft tokens. LogitSpec does not require an additional draft model and is a fully plug-and-play method, which can be easily implemented and integrated into existing LLM frameworks. Extensive experiments demonstrate that LogitSpec can effectively improve the retrieval performance, leading to a $1.8\times\sim 2.6\times$ speedup across all the evaluation benchmarks.
7 Limitations and Future Work.
We consider a few limitations and future works as follows. (i) While our LogitSpec is a fully plug-and-play SD framework, its real-world inference acceleration is less competitive. Our future works involve integrating LogitSpec into existing draft-model-based SD methods for further acceleration. (ii) Currently, LogitSpec retrieves relevant reference from the prompt, which may incur lower speedup when the prompt is short. We consider integrating an external database to boost the retrieval model as a future work. (iii) Our acceleration estimates rely on offline decoding with curated prompts and NVIDIA A100 GPUs, so the reported speedups may vary on other workloads or hardware. LogitSpec currently retrieves draft candidates from the on-the-fly prompt context only; short prompts or out-of-domain corpora reduce its benefits.
8 Ethic Statements
Inheritance of Model Behaviors.
LogitSpec is an inference acceleration framework designed to speed up existing Large Language Models without modifying their weights. As a speculative decoding method, it aims to losslessly recover the distribution of the target model. Consequently, LogitSpec inherits the ethical properties, biases, and potential safety risks of the underlying target LLM and draft model. It does not introduce new capabilities for generating harmful content, nor does it mitigate existing biases in the base models. Users should continue to apply standard safety guardrails and alignment techniques to the target models deployed with LogitSpec.
@@ Appendix A More Discussions to Related Work
We provide additional discussion of existing works in both draft-model-based speculative decoding and draft-model-free speculative decoding. An intuitive comparison is shown in Table 4.
@@
Besides the discussions in Section 2, we discuss additional works on draft-model-based speculative decoding. For example, HASS (Zhang et al., 2025) identifies the inconsistency between EAGLE’s training and inference phases, and proposes a multi-step training framework to address this. CORAL (Weng et al., 2025) proposes a cross-step representation alignment to address this problem. Judge Decoding (Bachmann et al., 2025) recognizes the potential of accepting high-quality but refused draft tokens to further improve the acceleration. Gumiho (Li et al., 2025a) demonstrates that the initial draft token is more important and proposes a hybrid model to combine serial and parallel draft heads.
Besides these methods that focus on the training process, the process of verification also draws extensive interest, mainly focusing on adaptive draft length. SpecDec++ (Huang et al., 2025) formulates the verification process as a Markov decision process to adaptively determine the draft length. OPT-Tree (Wang et al., 2024) proposes a method to search for the optimal tree structure that maximizes the mathematical expectation of the acceptance length in each decoding step. AdaEAGLE (Zhang et al., 2024b) proposes a novel framework to explicitly model the draft tree structure for EAGLE series models. PEARL (Liu et al., 2025) pioneers this direction by serving the draft model and the target model in parallel to achieve a segmented adaptive draft length.
@@
For layer sparsity, SPEED (Hooper et al., 2024) proposes a method to speculatively execute multiple future tokens in parallel with the current token using predicted values based on early-layer hidden states. FREE (Bae et al., 2023) proposes a shallow-deep module and a synchronized parallel decoding to improve the efficiency. EESD (Liu et al., 2024) proposes an early-exiting framework with a self-distillation method and leverages Thompson Sampling to regulate the generation processes. For retrieval-based SD, Token Recycling (Luo et al., 2024) proposes a method to utilize the dropped draft tokens and generate draft tokens via an adjacency matrix. Different from our method, Token Recycling is a query-dependent method that utilizes information from other queries, which may result in limitations when applied to real-world applications with complex and dynamic user inputs. SAM Decoding (Hu et al., 2024) utilizes a common text corpus and dynamic text sequence as retrieved sources and proposes a suffix automaton to efficiently obtain more accurate match positions.
Appendix B Details of Motivation Experiments
To investigate the prediction ability of the last logit, we conduct two motivation experiments to demonstrate its effectiveness and robustness. For the effectiveness of the last logit, as shown in Figure 2, we conduct autoregressive inference for the 6 models and record the logits for each decoded token. Then, for each decoded token $x_{i}$ which is sampled from the logit $p_{i-1}$, we investigate the rank of $x_{i}$ in $p_{i-2}$, i.e., the corresponding last logit, and visualize the statistics of the rank in Figure 2. For the robustness of the last logit, as shown in Figure 3, we conduct last logit decoding and investigate its mean accepted tokens per decoding step. Both experiments are conducted on a small subset of Spec-Bench, where we randomly sample 2 questions for each sub-category of MT, and 10 questions for other categories (Trans, Sum, QA, Math, RAG).
Why Not Use Next-Next-Next Tokens (NNNT)?
A natural question is whether we can extend the approach to predict even further tokens. We investigated this: the ground-truth NNNT appears in the top-60 last logits with $\geq$40% probability, showing the phenomenon does extend.
However, we did not use NNNT due to combinatorial explosion. If we speculate top-60 for both NNT and NNNT, the draft tree would contain at least $60^{2}=3{,}600$ tokens, far exceeding efficient verification capacity. As shown in Table 15 (Appendix E.8), tree capacity beyond 64-128 tokens yields diminishing returns as verification overhead dominates.
Our chosen approach (NNT + retrieval with K=64) represents the optimal efficiency-accuracy trade-off given current hardware constraints.
Intuition: Why Can Last Logits Predict Next-Next Tokens?
While our contribution is an empirical discovery, we offer intuition for why this phenomenon exists:
The last logit $p_{i-1}=P(x_{i}|x_{1},\ldots,x_{i-1})$ is computed from hidden states encoding the full context up to position i1i-1. While the model is trained to predict $x_{i}$, the rich contextual representation necessarily contains information about likely continuations beyond $x_{i}$—otherwise, the model could not generate coherent multi-token sequences.
Our findings suggest that top-ranked tokens in $p_{i-1}$ correlate with likely values of $x_{i+1}$, particularly in structured text where multi-token patterns (phrases, named entities, common collocations) are prevalent. This predictive signal exists without fine-tuning, making it immediately exploitable for inference acceleration.
We leave formal theoretical analysis of this phenomenon to future work.